# Quantum-inspired Techniques in Tensor Networks for Industrial Contexts

### v1

@@ 4 Use case studies
In this last section we will present the main use cases in which we can see a direct and significant potential of tensor networks techniques to industrial and business contexts.
@@ 4.2.1 Drug discovery
Drug simulation is a monumental challenge in contemporary science due to the complexity of molecular interactions and the enormous number of possible configurations. In this context, quantum computing offers a clear advantage as it is able to model and treat phenomena such as the electronic structure of molecules and interactions between atoms more adequately than classical computing. As explained above, the ability of tensor networks to efficiently represent quantum states may allow us to simulate certain molecules that are currently not possible with quantum computers.
Similarly, we can save time and costs of new drug discovery by predicting new therapeutic properties associated with new drugs. This can be achieved in several ways, one of them being the tensor decomposition [50]. To this end, we create a drug-gene-disease tensor, which accounts for the relationships found between these drug-gene, drug-disease and gene-disease pairs. With this, we formulate a tensor completion problem. The initial tensor will have only the direct relationships of the pairs we have found, while the rest of the more complex relationships will be the ones we have to determine. For this, we will perform a generalized tensor decomposition (GTD), to model part of these relationships, and we will apply a multilayer perceptron (MLP) to capture the rest of the relationships.
@@ 4.2.2 Medical image analysis
The analysis of images, and especially medical images, is a very complicated task, either because of the characteristics of these images, the noise in them or the high dimensionality that the problem can reach. However, as we have seen, tensor networks are ideal for cases with high dimensionality and noise.
In [51] authors present methods for classifying medical images, using exactly the scheme we explained in Ssec. 3.3, interpreting the images as an input vector and a feature map cosenoidal, and the MPO layer as a classifier. Thus, by contracting the tensor network we will obtain a vector that will tell us the probabilities of that image belonging to one group or another. This is done in combination with other more advanced machine learning techniques and finally we obtain an extremely high level of compression with levels of area under the ROC (Receiver Operating Characteristic) curve similar or superior to the methods with which it is compared. All this with much lower GPU requirements.
@@ 4.3 Simulation of quantum materials and topological materials
To deal with this, concrete ansatz are proposed [53, 54, 55, 56] which accurately and efficiently represent the minimum energy states of these systems. We start with an initial state and perform various transformations, such as DMRG or imaginary time evolution, to bring it down to its minimum energy state. Once this state is obtained, we can easily obtain properties of the state by applying the tensors of the quantities to be measured.
4.4 Battery simulation
Battery simulation is a very useful problem, but it also requires the study of quantum many-body systems. For this we can also use ansatz in tensor networks [57, 58], with which we can obtain properties of the minimum energy states.
@@ 4.5.1 Route Optimization
The main limitation of this method is the exponential scaling it suffers with the number of nodes due to the constraints, but iterative methods such as the one explored in [47] are able to limit this scaling successfully.
@@ 4.5.2 Post assignment optimization
A particular case of generalized TSP is the assignment of jobs to workers. Assume a set of workers who are already assigned to a set of jobs and we have a set of vacancies. The objective will be to redistribute workers to the set of jobs, including filled and vacant jobs, such that we maximize a certain cost function and there are only 0 or 1 workers per job. This problem has been studied and treated at the end of [45], using the formulation of [59], which seeks to maximize a combination between the productivity of the job and the compatibility of the worker assigned to it.
@@ 4.5.4 Machine task assignment optimization
This problem is addressed in [47], where a TN method with imaginary time evolution and constraint layers like those of [45] is combined with a genetic algorithm and an iterative method to reduce the time and memory scaling of such algorithm, taking advantage of properties of this type of problems. The main limitation of the method is that, although it has a high success rate, for certain cases it cannot obtain satisfactory solutions for a given memory limit.
@@ 4.6 Big Data
The work [40] presents different methods to deal with big data problems using tensor network decompositions, as well as other topics. On the other hand, in [41] a tensor network method for secret sharing in big data is presented, based on the decomposition into an MPS or a Tucker decomposition [60] of the original data tensor by means of consecutive TSVDs. Between each TSVD a tensor perturbation will be performed, by multiplying the new node obtained from the SVD by a $\Delta$ matrix, and the tensor to be decomposed in the next step by $\Delta^{-1}$. This achieves that the global tensor is the same, but the nodes of the representation are different. Finally, the first node of the representation is randomized. In this way, parts of the dataset can be stored in a distributed way while maintaining the security of the data, since no one will be able to obtain any information from the data until all the nodes of the representation are available. At the same time, it presents a way to perform distributed computing for big data.
@@ 4.8 Artificial Intelligence
As we saw in Ssec. 3.2, various models dedicated to artificial intelligence can be compressed using tensor network representations. Combining that with what was shown in Ssec. 3.3, we can deal with much larger and more complex problems. In addition, we can add various quantum machine learning techniques [4] to design new models with the features we need.
On the other hand, creating more compact models that can be inserted and trained directly on more modest devices, such as mobile devices, can allow experimentation with more models simultaneously and new applications.
We can also realize privacy-preserving machine learning techniques using tensor networks [63] techniques. This is of vital importance when working with sensitive data, such as medical records.
@@ 4.10 Anomaly detection
The work [39] discuss a supervised method precisely to be able to deal with this problem. By applying an ansatz product like the one we explained in Ssec. 3.3, and an operation in variational MPO that we will train, we can project the normal data to a hypersphere. In this way, the normal points will remain on the surface of the sphere, while the anomalous ones will go to the center or to the outside of the sphere. Moreover, the process can be performed very efficiently due to the structure of the tensor network to be contracted, scaling remarkably well with the size of the problem.
@@ 5 Conclusions
We have seen that tensor networks are a set of techniques with great applicability to the industrial context in an immediate way, with proven advantages and an extensive literature on the subject. We have explained the main features that make them especially attractive for high dimensionality problems and machine learning. We have also highlighted several applied cases and some of the techniques that have been used in the state of the art to address them, as well as their advantages and disadvantages.
Acknowledgement
The research leading to this paper has received funding from the Q4Real project (Quantum Computing for Real Industries), HAZITEK 2022, no. ZE-2022/00033.

### v2

@@ 4 Use case studies
In this last section we present use cases in which tensor-network techniques may have direct industrial applicability.
@@ 4.1.2 Interpretable predictions
4.1.3 Option pricing and risk simulations
Tensor trains, MPS, and QTT representations are increasingly used for option pricing, Greeks, and large-scale risk revaluation. The motivation is that full-grid PDE or tree-based methods suffer from dimensionality, while Monte Carlo methods may require many samples for stable Greeks or tail-risk estimates. Tensor-network methods can compress price surfaces, payoff tensors, operators, or parameter dependence, often through TT-cross or QTT representations [72, 6, 60, 61, 22].
The practical caveat is that the gains depend strongly on smoothness, parameter dimension, rank growth, calibration cost, and on whether the underlying pricing engine can be queried efficiently. In industrial settings, these methods should be compared against strong Monte Carlo, sparse-grid, Fourier, and reduced-order baselines for the target asset class and accuracy regime.
@@ 4.2.1 Drug discovery
Drug simulation is a monumental challenge in contemporary science due to the complexity of molecular interactions and the enormous number of possible configurations. Fault-tolerant quantum computing may eventually provide advantages for selected electronic-structure and quantum-chemistry tasks, but in the near term tensor-network methods such as DMRG and related ansatz-based techniques remain classical tools for exploiting structure in correlated molecular systems. Their usefulness depends on orbital ordering, entanglement structure, active-space selection, and the ranks required for the desired accuracy.
Similarly, we can save time and costs in drug discovery by predicting new therapeutic properties associated with new drugs. This is a data-driven tensor-decomposition and tensor-completion use case rather than a quantum many-body simulation [34]. To this end, we create a drug-gene-disease tensor, which accounts for the relationships found between these drug-gene, drug-disease and gene-disease pairs. With this, we formulate a tensor completion problem. The initial tensor will have only the direct relationships of the pairs we have found, while the rest of the more complex relationships will be the ones we have to determine. For this, we will perform a generalized tensor decomposition (GTD), to model part of these relationships, and we will apply a multilayer perceptron (MLP) to capture the rest of the relationships.
@@ 4.2.2 Medical image analysis
The analysis of images, and especially medical images, is a very complicated task, either because of the characteristics of these images, the noise in them or the high dimensionality that the problem can reach. Tensor networks can be useful in high-dimensional and noisy imaging tasks when the chosen tensorization exposes low-rank, local, or multiscale structure, and when truncation removes noise without discarding clinically relevant signal.
In [65] authors present methods for classifying medical images, using exactly the scheme we explained in Section 3.3, interpreting the images as an input vector, a trigonometric feature map, and the MPO layer as a classifier. Thus, by contracting the tensor network we will obtain a vector that will tell us the probabilities of that image belonging to one group or another. This is done in combination with other more advanced machine learning techniques and finally we obtain a high level of compression with levels of area under the ROC (Receiver Operating Characteristic) curve similar or superior to the methods with which it is compared. All this with much lower GPU requirements.
@@ 4.3 Simulation of quantum materials and topological materials
To deal with this, concrete ansatz are proposed [66, 63, 62, 42] which accurately and efficiently represent the minimum energy states of these systems. We start with an initial state and perform various transformations, such as DMRG in one-dimensional/MPS settings or related variational and imaginary-time algorithms for other tensor-network geometries, to bring it down to its minimum energy state. Once this state is obtained, we can easily obtain properties of the state by applying the tensors of the quantities to be measured.
4.4 Battery modelling and system identification
Battery modelling is an industrially relevant application where tensor methods can appear not only through quantum materials simulation, but also through data-driven system identification. The works considered here use tensor-network-based Volterra models for lithium-ion batteries, including MIMO settings, to represent nonlinear dynamical input-output relationships in compressed form [28, 29]. In this setting, the tensor network is not an ansatz for a many-body ground state; it is a structured representation of a high-order nonlinear model. The main industrial value is compact modelling of battery dynamics for estimation, control, and diagnostics, while the limitations are the availability and representativeness of data, the stability of the identified model, and the rank growth required to capture nonlinear behaviour.
@@ 4.5.1 Route Optimization
A complementary 2026 direction is to use tensor-network generative models for TSP. In TN-GEO, an MPS Born machine is trained over valid tours using permutation-based variables and autoregressive masking, so that generated samples are feasible by construction rather than obtained through an $N^{2}$ binary encoding with penalty terms [59]. This is a heuristic sampling strategy rather than a worst-case exact solver, but it illustrates an industrially relevant shift from penalty-heavy encodings to feasibility-aware sampling.
The main limitation of these methods is the exponential scaling induced by the constraints in the general case. Iterative or hybrid procedures such as the one explored in [4] can mitigate this scaling on the tested structured instances, but no general worst-case polynomial guarantee is implied.
@@ 4.5.2 Post assignment optimization
A particular case of generalized TSP is the assignment of jobs to workers. Assume a set of workers who are already assigned to a set of jobs and we have a set of vacancies. The objective will be to redistribute workers to the set of jobs, including filled and vacant jobs, such that we maximize a certain cost function and there are only 0 or 1 workers per job. This problem has been studied and treated at the end of [3], using the formulation of [16], which seeks to maximize a combination between the productivity of the job and the compatibility of the worker assigned to it. As in the route-optimization case, the practical performance depends on the structure of the instance and on the ranks required by the encoding.
@@ 4.5.4 Machine task assignment optimization
This problem is addressed in [4], where a TN method with imaginary time evolution and constraint layers like those of [3] is combined with a genetic algorithm and an iterative method to reduce the time and memory scaling of such algorithm, taking advantage of properties of this type of problems. The main limitation of the method is that, although it has a high success rate on the tested structured instances, for certain cases it cannot obtain satisfactory solutions for a given memory limit, and no general worst-case polynomial guarantee follows.
@@ 4.6 Big Data
The work [12] presents different methods to deal with big data problems using tensor network decompositions, as well as other topics. On the other hand, in [49] a tensor network method for secret sharing in big data is presented, based on the decomposition into an MPS or a Tucker decomposition [27] of the original data tensor by means of consecutive TSVDs. Between each TSVD a tensor perturbation will be performed, by multiplying the new node obtained from the SVD by a $\Delta$ matrix, and the tensor to be decomposed in the next step by $\Delta^{-1}$. This achieves that the global tensor is the same, but the nodes of the representation are different. Finally, the first node of the representation is randomized. In this way, parts of the dataset can be stored in a distributed way while restricting reconstruction of the original data under the assumptions of the specific secret-sharing protocol. A claim of perfect secrecy would require an explicit adversarial model and a proof, for example in terms of statistical independence or zero mutual information between unauthorized shares and the secret. At the same time, it presents a way to perform distributed computing for big data.
@@ 4.8 Artificial Intelligence
As discussed in Section 3.2, tensor networks can compress or parameterize components of artificial-intelligence models, especially when the chosen factorization aligns with the structure of linear layers or attention blocks. For industry, this is relevant for edge or on-premise deployment, privacy-sensitive environments, and reduced memory footprints, not only for parameter counting [70, 71, 41, 31].
Any practical evaluation should therefore compare tensor-network compression or tensorized adapters against quantization, pruning, distillation, LoRA/PEFT, and conventional low-rank adapters, reporting accuracy, compression ratio, latency, throughput, energy use, and memory rather than parameter count alone.
We can also realize privacy-preserving machine learning techniques using tensor networks [56]. This is of vital importance when working with sensitive data, such as medical records, but the same deployment-level benchmarking requirements apply.
4.9 Scientific computing, PDEs, and fluid dynamics
Scientific computing is an industrially important area in which tensor networks are increasingly used as compressed representations of operators, solution fields, flow maps, and parameter-dependent objects. Recent examples include MPS/MPO approximations of PDE flow maps, tensor-network fractional-step solvers for incompressible flow, and TT representations of high-dimensional combustion manifolds for CFD [15, 73, 17].
The attraction is clear: high-dimensional operators and lookup manifolds can sometimes be manipulated in compressed form rather than on full grids. The practical limitation is equally clear: performance depends on smoothness, separability, variable ordering, low rank, boundary and geometry handling, numerical stability, and comparison against strong classical reduced-order and sparse-grid baselines.
@@ 4.11 Anomaly detection
The work [77] discusses a supervised method precisely to be able to deal with this problem. By applying an ansatz product like the one we explained in Section 3.3, and an operation in variational MPO that we will train, we can project the normal data to a hypersphere. In this way, the normal points will remain on the surface of the sphere, while the anomalous ones will go to the center or to the outside of the sphere. Moreover, the process can be performed very efficiently due to the structure of the tensor network to be contracted, scaling remarkably well with the size of the problem.
@@ 5 Conclusions
We have seen that tensor networks are a versatile set of techniques for industrial contexts, with a growing literature and several promising application areas. We have explained the main features that make them attractive for structured high-dimensional problems and machine learning. We have also highlighted several applied cases and some of the techniques that have been used in the state of the art to address them, together with their advantages and limitations.
From an industrial perspective, tensor-network methods are most promising when the data, cost function or model exhibits locality, hierarchical structure, low effective entanglement, or correlations that can be aligned with the chosen network topology. They are less suitable when correlations are dense and global, when constraints induce high-rank projectors, or when exact guarantees are required for worst-case NP-hard families.
As of 2026, the industrial relevance of tensor networks is strongest in structured high-dimensional settings: quantum simulation and benchmarking, compressed machine-learning models, constrained optimization with exploitable structure, financial pricing and risk surfaces, scientific computing, and data-driven system identification. The main practical bottlenecks are tensorization choices, rank growth, contraction width, optimization stability, and hardware and software support. Consequently, industrial adoption should proceed through benchmarked pilots against strong classical baselines rather than through generic claims of exponential compression.